Solve by substitution.
x + 3y = 2
-x + y = 1
Thank you for helping!!

Answer by checkley75(3666)

x + 3y = 2 or x=2-3y
-x + y = 1
-(2-3y)+y=1
-2+3y+y=1
4y=1+2
4y=3
y=3/4 ans.
-x+3/4=1
-x=1-3/4
-x=1/4
x=-1/4 ans.
-1/4+3*3/4=2
-1/4+9/4=2
(-1+9)/4=2
8/4=2
2=2